# A new approach to $P-V$ phase transitions: Einstein gravity and holographic dark energy

### v1

@@ Abstract
In the framework of Einstein gravity, we study the thermodynamic equation state, $P=P(V,T)$, associated with a flat Friedmann-Lemaitre-Robertson-Walker (FLRW) universe with two dominant non-interacting fluids at late times related by the cosmological coincidence parameter. We find that under these circumstances, first-order $P-V$ phase transitions are admitted if the dark energy density and the cosmological coincidence parameter can be related to the geometry of the apparent horizon.
@@ I Introduction
Black holes have proven to be a very important environment to understand the linkage between General Relativity and thermodynamics during several decades. This has led to important results such as relating macroscopic properties associated to thermodynamics systems with the geometry of black holes and specially through the event horizon bekenstein ; haw ; haw2 ; dolan ; haykoda . However, it seems that this intimate connection emerges from Einstein equations since they can be interpreted as the first law of thermodynamics jacobson . In consequence, it is natural to think that the obtained results from the black holes dynamics can be translated into the cosmological scenario by replacing the usual fixed horizon by a dynamical object; in this latter construction such horizon corresponds to the so-called apparent horizon helo0 . This provides a natural framework to realize the universe as a thermodynamical system, see for instance references given in bekenstein2 as a first approach to this idea. It is worth mentioning that the nature of the cosmological horizon dictates the behavior of the thermodynamic quantities associated to it, i.e., there are subtle differences between the thermodynamics of a contracting cosmology and an expanding one, we refer the reader to Refs. helou ; posit , where these important distinctions between cosmologies were established, the difference in the thermodynamics emerging from black holes and white holes was also studied. From now on, we restrict ourselves to the case in which the cosmological principle prevails and the matter content of the universe is trapped by the apparent horizon surface in an expanding evolution.
Based on the above discussion, Einstein theory represents a robust scenario to explore the thermodynamics nature of the observable universe. Contrary to what is generally found in the literature under the thermodynamics system interpretation for an FLRW universe, see for instance modified see also various , where modified theories of gravity are considered; in this work we show that phase transitions in Einstein gravity are allowed by only assuming more than one fluid as matter sector. We consider the components of the dark sector as the dominant content of the universe and relate them by means of the cosmological coincidence parameter; this is the first and main consideration of our formulation. The dark energy and dark matter fluids are necessary in order to induce changes in the usual equation of state of the FLRW universe, namely
@@
However, adopting the standard thermodynamics description callen we find that the existence of critical points (and phase transitions) is determined by some conditions that are sensitive only under modifications that can be expressed in terms of the geometry of the apparent horizon, specifically its radius, $R_{A}$, as we will see later. At first glance this closure condition enforces dark energy to be of holographic type in order to write also the cosmological coincidence parameter in terms of $R_{A}$. Despite the limited this may seem, we have at hand several cases of interest that lead to phase transitions since we also provide some dark energy models physically motivated beyond the holographic scheme that fulfills the condition mentioned before. The approach presented here does not intend to oppose the idea of necessity of modified gravity scenarios to study the thermodynamics of the universe but to show that the Einstein theory is capable to exhibit a rich structure from the thermodynamics point of view with a simple refinement in the description of the matter sector.
The organization of the work is given as follows: in Section II, we describe the apparent horizon of an FLRW universe. We establish some relevant quantities associated with its thermodynamics description, such as surface gravity, temperature, and the unified first law. This is useful for relating the work density to the thermodynamic pressure, $P$. In Section II.1, we specialize the apparent horizon results to the case $k=0$ and use the standard procedure to establish the existence of $P-V$ phase transitions from the equation state of the universe, $P(V,T)$. We apply the method to the case in which the Friedmann equations describe two non-interacting fluids that are only related by the cosmological coincidence parameter. We describe the nature of the phase transition by exploring the behavior of the Gibbs free energy. The final comments of the work are given in Section III. $G=c=1$ units will be used throughout the work.
@@ II FLRW universe and apparent horizon
This section will briefly revise the thermodynamic description associated with the apparent horizon in an FLRW universe helo0 , such description is borrowed from the black hole dynamics haykoda . In our formulation we take into account two non-interacting cosmological fluids. The FLRW line element is usually given as
@@
Once we evaluate at the apparent horizon the above results, we can interpret $\Psi_{a}$ as the total energy flow at the horizon, and $W$ can be viewed as the work done by the change of the apparent horizon. This quantity can be related to the thermodynamic pressure, $P$, since can be considered as the conjugate variable of the thermodynamic volume from the first law of thermodynamics, we have $W:=P$.
@@ II.1 $P-v$ phase transitions
It is worth mentioning that the temperature depends on the $R_{A}$ and its change along cosmic evolution, the case $\dot{R}_{A}/(2HR_{A})\ll 1$ in the latter expression resembles the temperature of a spherically symmetric black hole with horizon radius $R_{A}$. In general, the temperature on the horizon is defined by the characteristics of the cosmic expansion generated by the energy content of the universe through the deceleration parameter $q$, which is defined as $1+q:=-\dot{H}/H^{2}$. This expression is useful to notice that the positivity of the temperature is maintained for quintessence and phantom dark energy.
@@
where a barotropic equation of state of the form, $p=\omega\rho$, was considered for each specie and $\omega_{\mathrm{m}}=0$ to describe a pressureless fluid for the matter sector. Besides, the Eq. (16) was used to substitute $\dot{R}_{A}$ by $T_{A}$. Let’s consider the usual definition for the specific volume, $v=2R_{A}$. The latter expression is the thermodynamic equation of state for the FLRW universe in which we consider two non-interacting fluids to describe the universe’s matter content. The conditions to have a $P-v$ phase transition in the system are
@@
and when applied to (21), no critical points are found; therefore, the model does not admit $P-v$ phase transitions. Notice that the lack of critical points keeps if the form of the coincidence parameter is written in terms of the redshift or scale factor. Then, the coincidence parameter must be related to the specific volume $v$ (or $R_{A}$); this is crucial to allow $P-v$ phase transitions in Einstein gravity. According to Eq. (19), the demanding condition discussed above on $r$ to allow phase transitions in our scheme restricts severely the form of the energy density associated with the dark energy sector. However, if we adopt the usual holographic dark energy formula as warming up case, we can write directly the desired expression, $r=r(R_{A})$. For generalization purposes, we propose the following form for the energy density of the dark energy sector
@@
where $\beta$ and $n$ are real parameters with $n\geq 1$ being an integer. In the context of holographic dark energy, $\rho_{\mathrm{de}}\propto 2L^{-2}$, where the characteristic length can be given as the inverse of the Hubble scale, $H$. Inserting Eq. (23) in (19) we can write the following expression
@@
Notice that for $n=1$ the second term in the last expression is constant, therefore no critical points can be found in this case since resembles the usual equation of state (1). In terms of the specific volume defined above, for $n>1$ we obtain from the conditions (22), the following critical point
@@
In the following, we will focus on some specific values for $n$ that are physically justified within the cosmological context. The value $n=2$ corresponds to a dark energy model recently proposed in Refs. transient1 ; transient2 and exhibits an early phantom scenario with de Sitter final stage at the far future of the cosmological evolution. Beyond Einstein’s gravity, it is worth mentioning that a $H^{4}$ term also appears in the framework of cubic gravity as a correction term for cosmic evolution cubic . The case $n=3$ will also be considered, a generalization of cubic gravity admits correction terms of the form $H^{6}$ in the acceleration equation h6 ; sixth . Finally, the cosmological evolution of quartic gravity was explored in Ref. oliva and corrected by a term of the form $H^{8}$, corresponding to the case $n=4$. However, powers of the Hubble parameter in the Friedmann equations are usually found in various theories of gravity including Lovelock and scalar-tensor theories, see for instance Refs. given in various . Using the equations (26), (27) and (28) we can write
@@
which is a constant value and independent of the parameter $\beta$. For $n=1$ the latter expression diverges, therefore $P-v$ phase transitions are expected for $n>1$. On the other hand, if we specialize the above expression for the values $n=2,3,4$ we obtain $3/8,(10/9)(3/8),(7/6)(3/8)$, respectively. Notice that the constant obtained in each case is proportional to the universal ratio obtained from the van der Waals theory given by $3/8$ which establishes an upper bound for real gases. For the fluids discussed here this value represents a lower bound for the ratio $(P_{c}v_{c})/T_{c}$. This indicates that for $3/8$ the existence of phase transitions in these cosmological fluids is consistent with the mean-field theory description and in the cases where deviations from $3/8$ are obtained, the phase transitions can be due to interactions between the constituents of the fluid beyond the van der Waals gas description vanderwaals .
In order to have dimensionless quantities, we define the following reduced variables: $\bar{P}:=P/P_{c}$, $\bar{T}:=T/T_{c}$ and $\bar{v}:=v/v_{c}$, thus Eq. (25) can be written as follows
@@
Using the plot for the reduced Gibbs free energy against the reduced pressure, $\bar{P}$, we can characterize the phase transition observed in the $P-v$ diagram given above. We depict the reduced Gibbs free energy behavior in Fig. (2). We obtain $\bar{G}$ from the standard thermodynamic relation
@@
Notice that the size of the loop characterizing the first-order phase transition decreases for two different situations: (i) different dark energy models as the value of $n$ increases and (ii) for a fixed $n$ but as we increase the value of the temperature. The first case indicates that for higher powers in $H$, the dark energy models proposed here are more stable thermodynamically speaking.
Based on the thermodynamics construction discussed above, we extend our formalism to dark energy models, which contain a linear term in the first derivative of the Hubble parameter. We propose the following modification for the energy density (23)
@@
with $\beta$ and $\gamma$ being positive arbitrary parameters; for $n=1$, the Granda-Oliveros holographic model is recovered grandaoliveros , which in turn results to be an extension of the Ricci holographic model ricci . Notice that a generalization of this type is permitted since the first derivative of $H$ is related to the temperature of the horizon by means of Eq. (16). The inclusion of the first derivative of the Hubble parameter into the dark energy density can also be motivated with scenarios beyond the holographic approach. In the context of QCD, minimal extensions of the ghost dark energy take into account the first derivative of $H$. In this scenario, the magnitude of the vacuum energy is corrected by the contribution of some ghost fields given as $\rho\propto H\Lambda^{3}_{QCD}$, being $\Lambda_{QCD}$ the QCD mass scale. See, for instance, the references given in qcd for a complete guide on the topic and the cases studied when the first derivative of $H$ is considered as a correction term to the ghost dark energy.

### v2

# A new approach to $P-V$ phase transitions: Einstein gravity and holographic type dark energy

@@ Abstract
In the framework of Einstein’s gravity, we study the thermodynamic equation state, $P=P(V,T)$, associated with a flat Friedmann-Lemaitre-Robertson-Walker (FLRW) universe. In this scenario, we consider the components of the dark sector as non-interacting fluids that dominate the universe’s energy content at late times. Under these circumstances, the functional structure of the cosmological coincidence parameter plays a relevant role in admitting first-order $P-V$ phase transitions; specifically, the dark energy density and the coincidence parameter must be given in terms of the radius of the apparent horizon.
@@ I Introduction
Black holes have been a very important environment for understanding the linkage between general relativity and thermodynamics for several decades. This has led to important results such as relating macroscopic properties associated with thermodynamic systems with the geometry of black holes, especially through the event horizon bekenstein ; haw ; haw2 ; dolan ; haykoda . However, this intimate connection seems to emerge from Einstein’s equations since they can be interpreted as the first law of thermodynamics jacobson . Consequently, it is natural to think that the results obtained from the black hole dynamics can be translated into the cosmological scenario by replacing the usual fixed horizon with a dynamical object; in this latter construction, such horizon corresponds to the so-called apparent horizon helo0 . This provides a natural framework to realize the universe as a thermodynamical system, see for instance references given in bekenstein2 as a first approach to this idea. It is worth mentioning that the nature of the cosmological horizon dictates the behavior of the thermodynamic quantities associated with it, i.e., there are subtle differences between the thermodynamic of a contracting cosmology and an expanding one, we refer the reader to Refs. helou ; posit , where these important distinctions between cosmologies were established, the difference in the thermodynamics emerging from black holes and white holes was also studied. From now on, we restrict ourselves to the case in which the cosmological principle prevails, and the universe’s matter content is trapped by the apparent horizon surface in an expanding evolution.
Based on the above discussion, Einstein’s theory represents a robust scenario to explore the thermodynamic nature of the observable universe. Contrary to what is generally found in the literature under the thermodynamic system interpretation for an FLRW universe, see for instance modified see also various , where modified theories of gravity are considered; in this work, we show that phase transitions in Einstein gravity are allowed by only assuming more than one fluid as matter sector. We consider the dark sector components as the universe’s dominant content and relate them using the cosmological coincidence parameter; this is the first and main consideration of our formulation. The dark energy and dark matter fluids are necessary to induce changes in the usual equation of the state of the FLRW universe, namely
@@
However, adopting the standard thermodynamic description callen , we find that the existence of critical points (and phase transitions) is determined by some conditions that are sensitive only under modifications that can be expressed in terms of the geometry of the apparent horizon, specifically its radius, $R_{A}$, as we will see later. At first glance, this closure condition enforces dark energy to be of the holographic type to write also the cosmological coincidence parameter in terms of $R_{A}$. Despite the limited this may seem, we have at hand several cases of interest that lead to phase transitions since we also provide some dark energy models physically motivated beyond the holographic scheme that fulfills the condition mentioned before. The approach presented here does not intend to oppose the idea of the necessity of modified gravity scenarios to study the thermodynamics of the universe but to show that Einstein’s theory is capable of exhibiting a rich structure from the thermodynamic point of view with a simple refinement in the description of the matter sector.
The organization of the work is given as follows: in Section II, we describe the apparent horizon of an FLRW universe. We establish relevant quantities associated with its thermodynamic description, such as surface gravity, temperature, and the unified first law. This is useful for relating the work density to the thermodynamic pressure, $P$. In Section II.1, we specialize the apparent horizon results to the case $k=0$ and use the standard procedure to establish the existence of $P-V$ phase transitions from the equation state of the universe, $P(V,T)$. We apply the method to the case where the Friedmann equations describe two non-interacting fluids only related by the cosmological coincidence parameter. We describe the nature of the phase transition by exploring the behavior of the Gibbs free energy. The final comments of the work are given in Section III. $G=c=1$ units will be used throughout the work.
@@ II FLRW universe and apparent horizon
This section will briefly revise the thermodynamic description associated with the apparent horizon in an FLRW universe helo0 ; such description is borrowed from the black hole dynamics haykoda . In our formulation, we take into account two non-interacting cosmological fluids. The FLRW line element is usually given as
@@
Once we evaluate at the apparent horizon the above results, we can interpret $\Psi_{a}$ as the total energy flow at the horizon, and $W$ can be viewed as the work done by the change of the apparent horizon. This quantity can be related to the thermodynamic pressure, $P$ since it can be considered as the conjugate variable of the thermodynamic volume from the first law of thermodynamics; we have $W:=P$.
@@ II.1 $P-v$ phase transitions
It is worth mentioning that the temperature depends on the $R_{A}$ and its change along cosmic evolution, the case $\dot{R}_{A}/(2HR_{A})\ll 1$ in the latter expression resembles the temperature of a spherically symmetric black hole with horizon radius $R_{A}$. In general, the temperature on the horizon is defined by the characteristics of the cosmic expansion generated by the energy content of the universe through the deceleration parameter $q$, which is defined as $1+q:=-\dot{H}/H^{2}$. This expression is useful to notice that the positive of the temperature is maintained for quintessence and phantom dark energy.
@@
where a barotropic equation of state of the form, $p=\omega\rho$, was considered for each specie and $\omega_{\mathrm{m}}=0$ to describe a pressureless fluid for the matter sector. Besides, the Eq. (16) was used to substitute $\dot{R}_{A}$ by $T_{A}$. Consider the usual definition for the specific volume, $v=2R_{A}$. The latter expression is the thermodynamic equation of state for the FLRW universe in which we consider two non-interacting fluids to describe the universe’s matter content. The conditions to have a $P-v$ phase transition in the system are
@@
and when applied to (21), no critical points are found; therefore, the model does not admit $P-v$ phase transitions. Notice that the lack of critical points in the coincidence parameter is written in terms of the redshift or scale factor. Then, the coincidence parameter must be related to the specific volume $v$ (or $R_{A}$); this is crucial to allow $P-v$ phase transitions in Einstein gravity. According to Eq. (19), the demanding condition discussed above on $r$ to allow phase transitions in our scheme severely restricts the form of the energy density associated with the dark energy sector. However, if we adopt the usual holographic dark energy formula, we can write directly the desired expression, $r=r(R_{A})$. This represents an interesting scenario since, as discussed in colgain , the holographic principle is a propitious arena to confront theory, model assumptions, and recent observations. The holographic principle is the most important cornerstone of quantum gravity and allows us to describe the physical quantities inside the universe, including the energy density of the dark energy sector, by some quantities on the universe’s boundary. Based on the dimensional analysis, only two quantities can be used to construct the dark energy density: the reduced Planck mass, $M_{P}$, and the characteristic length $L$ hp ; Nojiri:2005pu ; Nojiri:2019yzg
$\rho_{\mathrm{de}}=C_{1}M^{4}_{P}+C_{2}M^{2}_{P}L^{-2}+C_{3}L^{-4}+\mathcal{O}(L^{-6})+...$ (23)
where the parameters $C_{1},C_{2},C_{3}$ are constant. The $C_{1}$ term is not compatible with the holographic principle, it is well established that such a term leads to the cosmological constant problem, then the expansion given above should start from the second term [^4] It is important to stress out that the form of (23) comes from the application of the renormalization group method in cosmology, i.e., the full vacuum energy is a physical observable when the leading quantum contributions to it are characterized, such contributions are given by an even-power law of the Hubble expansion rate, $\rho_{\Lambda}(H)\propto\rho^{0}_{\Lambda}+M^{2}_{P}H^{2}+\mathcal{O}(H^{4})+\mathcal{O}(H^{6})+...$, where $\rho_{\Lambda}$ is the renormalized part of the complete QFT structure of the vacuum energy. For a complete discussion, see Ref. Nojiri:2017opc ; qft and references therein.. According to our scheme, the Hubble scale gives the most convenient election for the characteristic length, which is defined as $L:=1/H$. For generalization purposes, we used the spirit proposed by Nojiri:2017opc , where the authors introduced a generalized holographic dark energy model of the form
$L=L(L_{p},\dot{L}_{p},\ddot{L}_{p},...,L_{f},\dot{L}_{f},\ddot{L}_{f},...,H,\dot{H},\ddot{H},...,t_{s}).$ (24)
The infrared Nojiri-Odintsov cutoff corresponds to a combination of the FRW universe parameters: the Hubble parameter, particle ($L_{p}$) and future horizons ($L_{f}$), and their derivatives, cosmological constant and the universe lifetime, $t_{s}$, (if finite). In our new approach to finding and describing the $P-v$ phase transition, we use the simplest choice of the cutoff (24), which corresponds to $L=L(H)$ (and its generalization $L=L(H,\dot{H}$)) because it is possible to map such holographic dark energy to modified gravity or gravity with general fluid. Therefore, we consider the following form for the energy density of the dark energy sector
@@
where $\beta$ and $n$ are real parameters with $n\geq 1$ being an integer. The existence of possible analytic solutions for the critical points equations (22) remain in the form that we choose the holographic dark energy, Eq. (25) in our case. The dark energy model (25) can be related to the geometry of the apparent horizon through the specific volume $v$ (or $R_{A}$) in a simple way due to Eq. (14), in consequence the coincidence parameter acquires the desired functional form, $r=r(v)$; this is the critical ingredient to characterize in an analytic path the $P-v$ phase transitions in Einstein’s gravity, as commented above for the critical conditions (22), notice that the form given above for the energy density covers some of the terms given in the expansion (23), we will focus on some cases with physical meaning, as we will discuss below. The classical stability of the dark energy models emerging from Eq. 25 is discussed in the appendix A. Inserting Eq. (25) in (19) we can write the following expression
@@
Notice that for $n=1$, the second term in the last expression is constant. Therefore, no critical points can be found in this case since it resembles the usual equation of state (1). In terms of the specific volume defined above, for $n>1$ we obtain from the conditions (22), the following critical point
@@
In the following, we will focus on specific values for $n$ that are physically justified within the cosmological context. The value $n=2$ corresponds to a dark energy model recently proposed in Refs. transient1 ; transient2 with the following important properties: the dark energy parameter state can be written in terms of the coincidence parameter defined as, $r:=\rho_{\mathrm{m}}/\rho_{\mathrm{de}}$, yielding
$\omega_{\mathrm{de}}=-1-2(1+\omega_{\mathrm{m}})\frac{r(z)}{1-r(z)},$ (31)
where the coincidence parameter in terms of the redshift, $z$, has the explicit form
$r(z)=\frac{\lambda(1+z)^{3}}{\left(1+\sqrt{1-\lambda(1+z)^{3}}\right)^{2}},$ (32)
being $\lambda$ a constant parameter defined as, $\lambda:=4[1-\Omega_{\mathrm{m,0}}]\Omega_{\mathrm{m,0}}$, where $\Omega_{\mathrm{m,0}}$ is the fractional energy density associated to the matter sector at present time; then the model describes an early phantom scenario since at the beginning of the dark energy dominance epoch given at $r=1$, the parameter state $\omega_{\mathrm{de}}$ diverges. In the case of the future evolution of these dark energy models differs considerably despite the similarities between ghost dark energy and generalized holographic dark energy; the full classification for future type V singularities can be seen in Ref. deHaro:2023lbq ; Trivedi:2023zlf ; Trivedi:2022ngt . However, its energy density remains bounded, thus the singular behavior is obtained only in the parameter state (pressure) of the fluid. As can be seen in Ref. transient1 ; transient2 previously, our scheme admits a phantom scenario whose singularity lies only on the parameter state of dark energy and appears in the deceleration-acceleration transition stage; from there, the model tends to a future de Sitter evolution, i.e., our approach allows a transient early (not future) phantom regime. A signal of stability for this dark energy model is obtained from the positivity of the squared adiabatic sound speed. In fact, at the far future ($z=-1$) from Eq. (32) we can observe that $\omega_{\mathrm{de}}(z\rightarrow-1)\rightarrow-1$, therefore the model exhibits a future transition to de Sitter expansion, further details of the model are discussed in transient1 ; transient2 . Beyond Einstein’s gravity, it is worth mentioning that a $H^{4}$ term also appears in the framework of cubic gravity [^5] In this scenario, the Einstein-Hilbert action is corrected by the term $P$ as follows $S=\int d^{4}x\sqrt{-g}\left[\frac{1}{2\kappa}(R-2\Lambda)+\alpha P\right],$ (33) where $P$ is a general non-topological cubic term that involves curvature invariants. as a correction term for cosmic evolution cubic . The case $n=3$ will also be considered, a generalization of the type, $P\rightarrow f(P)$ in action (33), being $f$ an arbitrary function of the cubic term, admits correction terms of the form $H^{6}$ in the acceleration equation sixth , in this kind of model the parameter state of the dark energy sector also experiences a future transition from quintessence (phantom) to de Sitter evolution, the behavior of dark energy during the cosmological evolution, in this case, depends on the election of the function $f$. Finally, the cosmological evolution of quartic gravity was explored in Ref. oliva and corrected by a term of the form $H^{8}$, which corresponds to the case $n=4$; for this scenario, was found that a matter dominated epoch is smoothly connected to a late time acceleration described by a de Sitter phase. It is worth mentioning that gravitational theories discussed above represent curvature corrections to General Relativity whose vacuum spectrum consists of a graviton and is ghost-free cubic ; sixth ; oliva , a relevant characteristic of these modified scenarios is that contrary to what turns out to be a generic feature of higher curvature contributions, the Friedmann equations preserve their second order in the derivatives, this feature avoids the presence of ill behaviors as ghost or massive modes. Within the cosmological scenario, only powers of $H$ are expected to contribute to the critical density in order to maintain the aforementioned properties of the theory; for details, see cubic ; sixth ; oliva . Thus, we consider two principal arguments to construct our energy density for the dark energy sector: (i) the simplest form of the generalized holographic cutoff given in Eq. (24) and (ii) in order to avoid the propagation of extra degrees of freedom which is guaranteed by a set of second order Friedmann equations, makes dark energy as the one given in Eq. (25), a suitable election to contribute to the energy density of the universe.
Additionally, powers of the Hubble parameter in the Friedmann equations are usually found in various theories of gravity, including Lovelock and scalar-tensor theories, such as Refs. given in h6 ; various .
Using the equations (28), (29) and (30) we can write
@@
which is a constant value and independent of the parameter $\beta$. For $n=1$ the latter expression diverges, therefore $P-v$ phase transitions are expected for $n>1$. On the other hand, if we specialize the above expression for the values $n=2,3,4$ we obtain $3/8,(10/9)(3/8),(7/6)(3/8)$, respectively. Notice that the constant obtained in each case is proportional to the universal ratio obtained from the van der Waals theory given by $3/8$, which establishes an upper bound for real gases. For the fluids discussed here, this value represents a lower bound for the ratio $(P_{c}v_{c})/T_{c}$. This indicates that for $3/8$, the existence of phase transitions in these cosmological fluids is consistent with the mean-field theory description, and in the cases where deviations from $3/8$ are obtained, the phase transitions can be due to interactions between the constituents of the fluid beyond the van der Waals gas description vanderwaals .
To have dimensionless quantities, we define the following reduced variables: $\bar{P}:=P/P_{c}$, $\bar{T}:=T/T_{c}$ and $\bar{v}:=v/v_{c}$, thus Eq. (27) can be written as follows
@@
Using the plot for the reduced Gibbs free energy against the reduced pressure, $\bar{P}$, we can characterize the phase transition observed in the $P-v$ diagram given above. Fig. (2) depicts the reduced Gibbs free energy behavior. We obtain $\bar{G}$ from the standard thermodynamic relation
@@
Notice that the size of the loop characterizing the first-order phase transition decreases for two different situations: (i) different dark energy models as the value of $n$ increases and (ii) for a fixed $n$ but as we increase the value of the temperature. The first case indicates that for higher powers in $H$, the dark energy models proposed here are more stable thermodynamically speaking, and in consequence, the $P-v$ diagram in such cases will be more similar to the one associated with the ideal gas. Hence, no phase transition will be observed. Aside from the physical justification from diverse cosmological models, the thermodynamic description imposes important bounds on the values of the exponent $n$ that can be considered in the dark energy density (25) to obtain phase transitions in the model. As can be observed in Fig. (1), the coexistence phase per isotherm in the $P-v$ diagram is obtained for small values of the specific volume since we have $v\propto H^{-1}$, the phase transition takes place at the future. This means that phase transitions in the thermodynamic context reveal the transient behavior of dark energy and its passage from phantom (quintessence) to a future de Sitter expansion, as discussed before for the cosmological models considered here. It is of the utmost importance that thermodynamic phase transitions occur in the future evolution. If a radical phase transition had occurred in the recent past, it would have left observable imprints diaz .
Based on the above-mentioned thermodynamic construction and the generalized holographic dark energy models presented in Nojiri:2017opc , we extend our formalism developed to holographic dark energy models, which contain a linear term in the first derivative of the Hubble parameter. We use the following modification for the energy density, (25)
@@
with $\beta$ and $\gamma$ being positive arbitrary parameters; for $n=1$, the Granda-Oliveros holographic model is recovered grandaoliveros , which in turn results to be an extension of the Ricci holographic model ricci , see also ricci2 for an alternative extension of this model. Notice that a generalization of this type is permitted since the first derivative of $H$ is related to the temperature of the horizon through Eq. (16). The inclusion of the first derivative of the Hubble parameter into the dark energy density can also be motivated with scenarios beyond the holographic approach. In the context of QCD, minimal extensions of the ghost dark energy consider the first derivative of $H$. In this scenario, the magnitude of the vacuum energy is corrected by the contribution of some ghost fields given as $\rho\propto H\Lambda^{3}_{QCD}$, being $\Lambda_{QCD}$ the QCD mass scale. See, for instance, the references given in qcd for a complete guide on the topic and the cases studied when the first derivative of $H$ is considered as a correction term to the ghost dark energy.